OpenStudy (anonymous):

for the followoingf system of equations,what is the x-value of the solution? -x+2y=6, 6y+x+18

OpenStudy (mathstudent55):

Is the second equation 6y + x = 18 (= instead of +)? If so, then write the equation with the variables in the same order: x + 2y = 6 x + 6y = 18

OpenStudy (anonymous):

it is 6y=x+18

OpenStudy (mathstudent55):

Ok, then let's start again. x + 2y = 6 6y = x + 18 First, in the second equation, subtract x from both sides. x + 2y = 6 -x + 6y = 18

OpenStudy (mathstudent55):

Since you want to solve for x, let's eliminate y. To do so, you have to have the y-terms of both equations in a way so that when you add them together, they add up to zero. Since you have 6y in the 2nd equation and 2y in the first equation, what do you need to do to the 2y to turn it into -6y, so when you add it to 6y it cancels out?

OpenStudy (anonymous):

times by -3

OpenStudy (mathstudent55):

Exactly, so multiply both sides of the first eq by -3 and then write below that the second equation.

OpenStudy (anonymous):

-x+6y=18 -3x+-6y=6

OpenStudy (mathstudent55):

Remember to multiply every term of both sides.

OpenStudy (mathstudent55):

x + 2y = 6 Multiply both sides by -3: -3(x + 2y) = -3(6) -3x - 6y = ?

OpenStudy (anonymous):

-18

OpenStudy (mathstudent55):

Right. So now write the new first equation. then below that write the second equation.

OpenStudy (mathstudent55):

-3x - 6y = -18 -x + 6y = 18 ---------------

OpenStudy (mathstudent55):

That's what you should have. I wrote a line under it because now you need to add both equations together.

OpenStudy (anonymous):

not sure what u mean

OpenStudy (anonymous):

wouldnt it just be -4x

OpenStudy (mathstudent55):

You see the way I wrote the equations one abopve the other with like terms lined up. Just add like terms.

OpenStudy (mathstudent55):

-3x - 6y = -18 -x + 6y = 18 --------------- -4x + 0 = 0

OpenStudy (mathstudent55):

So if -4x = 0, now you divide both sides by -4, and you get x = ?

OpenStudy (anonymous):

o

OpenStudy (mathstudent55):

Right, x = 0 Since you were only asked for the x-value of the solution, you're done.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

you're welcome